OpenStudy (kimjafo):

Convert the equation to the standard form for a hyperbola by completing the square on x and y. y^2- 25x^2 + 4y + 50x - 46 = 0

OpenStudy (zzr0ck3r):

can you complete the square on \(-25x^2+50x-46\)?

OpenStudy (zzr0ck3r):

if not, can you find the vertex?

OpenStudy (kimjafo):

I tried before and got it wrong

OpenStudy (zzr0ck3r):

do you know how the vertex relates to completing the square?

OpenStudy (kimjafo):

No

OpenStudy (zzr0ck3r):

completing the square means to take a quadratic of the form \(ax^2+bx+c\) to the form \(a(x-h)^2+k\) where \((h,k)\) is the vertex. Do you know what a vertex is?

OpenStudy (kimjafo):

I believe I do

OpenStudy (zzr0ck3r):

ok forget the vertex thing, we will just complete the square a different way. Using the vertex makes it easier but if you are not used to working with the vertex then I assume your teacher wants you to do it another way.

OpenStudy (kimjafo):

I do an online schooling, so I teach myself of time

OpenStudy (zzr0ck3r):

\(y^2- 25x^2 + 4y + 50x - 46 = 0\) \(y^2+4y- 25x^2 + 50x = 46\) We need to complete the square twice for this problem. But one is harder than the other. \(y^2+4v-25(x^2-2)=46\) \(y^2+4y-25(x-1)^2=46-25(-1)^2\) \(y^2+4y-25(x-1)^2=21\) What I did was factor out the -25 from the x terms and then complete the square for the quadratic in x \((y+2)^2-25(x-1)^2=21+(-2)^2\) \((y+2)^2-25(x-1)^2=5^2\)

OpenStudy (zzr0ck3r):

make sure that you understand that I completed the square twice and then I will show you how to do it.

OpenStudy (kimjafo):

Ah! I was doing the problems wrong 😅😅 thanks!